OpenStudy (anonymous):

Solve each equation. Give all solutions in radians in the interval 0 < x < 2pi. - a) 2sin x + 3 = 2 b) cosx + 1= 0 c) 3tan^2x = 1 d) 2sin^2x + sinx = 1 e) tanx cosx - tanx = 0

OpenStudy (anonymous):

way too many questions. sorry.

OpenStudy (anonymous):

Give all solutions in radians in the interval \[0\le x < 2\pi \]

OpenStudy (anonymous):

a) \[2\sin x + 3 = 2\]

OpenStudy (anonymous):

2 sin x = -1 sin x = -1/2 x = 7pi/6 and 11pi/12

OpenStudy (anonymous):

You are probably wondering where I get 7pi/6 and 11pi/12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Watch.......we know that sin (30 degrees or pi/6) = 1/2 The sine is negative in quadrants 3 and 4. So if we want to know where sin x = -1/2..we know it will be - in quadrants 3 and 4. In quadrant 3, we just take 180 + 30 = 210 = 7pi/6 In quadrant 4, we do 360 - 30 = 330 = 11pi/12

OpenStudy (anonymous):

response??

OpenStudy (anonymous):

is it not 11pi /6

OpenStudy (anonymous):

sorry, you are correct, 330 degrees = 11pi/6. I stand corrected!

OpenStudy (anonymous):

7pi/6 and 11pi/6

OpenStudy (anonymous):

I tried to give you an explanation "short and sweet"...but there's an entire chapter in your text about this.

OpenStudy (anonymous):

so is all that my answer?

OpenStudy (anonymous):

there are 2 solutions for x, 7pi/6 and 11pi/6.

OpenStudy (anonymous):

7pi/6 and 11pi/6 that is my answer correct?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

b) cosx + 1 = 0

OpenStudy (anonymous):

@Easyaspi314